HUBBLE SPACE TELESCOPE DAILY REPORT #963 PERIOD COVERED: 0800 EST Daily Status Report as of 244/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/RD 4545 (Outgassing of Ceres - Cycle 3 High) The Faint Object Spectrograph was used to make ultraviolet spectral observations of the asteroid Ceres to attempt and confirm observations made with the IUE which detected OH emissions. The decision observation was received in real-time and displayed. It indicated that the target was perfectly centered. No pointing correction maneuver was necessary. The first science observations were received and displayed for investigators. A few guide star recentering events were reported during the observation set, but the observations were reported to look okay. 1.2 Completed WFC 4929 (WF/PC Cycle 2 : Improved Dark Calibration) The Planetary Camera was used to make two dark calibration images. The two exposures lasted 1200 seconds each, and the camera was pre-flashed before the observations. Both observations were reported to be normal exposures. 1.3 Completed WFC/PC 5026 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) The Wide Field Camera and the Planetary Camera were used to make two internal flat calibration images. These images will be used to look for measles and daisies (i.e., different forms of contamination), and to monitor scattered light. Both observations were executed as scheduled, and no problems were reported. 1.4 Completed FOC/96 4546 (Imaging of UV Bright Stars in the galactic Bulge) The Faint Object Camera was used to make observation of the bright ultraviolet star SGR-I which is located in Baade's Window. The observations were executed as scheduled, and no problems were reported. 1.5 Completed Three Sets of WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) The Wide Field Camera was used to make a total of two internal flat calibrations images. Both images were reported to be normal. 1.6 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.7 Completed PC 4712 (WF/PC Cycle 2 Calibration: Measles Monitor) The Planetary Camera was used to make Earth Calibration Flat images to monitor fo a form of contamination known as measles. The observations used filter F517N, both with and without filter F122M. The OSS monitored the observation and assessed the data quality. No problems were reported. 1.8 Completed WFC 5013 (WF/PC Cycle Flats Narrow. 2) The Wide Field Camera was used to make Earth flat calibration images to be used for flat fielding of WFC observations. The observations used the narrow band filters, both with and without neutral density filters. The observations were received and displayed. No problems were reported. 1.9 Completed PC 4714 (WF/PC Cycle Calibration: PC Internal Delta Flats) The Faint Object Camera was used to make one internal flat calibration image which will be used for the restoration of Earth flat images, which were altered due to changes in quantum efficiency which occurred after decontamination operations. The observations was received and displayed in the OSS. It was reported to look okay. 1.10 Completed FOS/BL 4662 (Exploring Triton's UV Spectrum: The First 1800-2500 AO UV Survey) The Faint Object Spectrograph was used to make observations for the moon Triton between 1800 and 2500 A0. This region has not been observed before. They are looking for CO, NO, and hydrocarbon absorption lines in its spectra. The OSS monitored the observations, and assessed the data quality. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 13 Successful : 13 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Significant HSTARs: None 2.4 Operations Note: The JIS #2 Simulation was initiated yesterday (Tuesday, Aug. 31, 1993 at 2200Z). The simulation will cover sequences 50 through 60 (PREP FOR SA RETRACT through SAFE MCU), and is scheduled to last for approximately 37 hours. So far all simulation objectives have been met. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Four Sets of PC 4714 (WF/PC Cycle 2 Calibration: PC Internal Delta Flats) 3.2 Complete Two Sets of PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2 ) 3.3 Complete FOC/96 4834 (Crane, Augmentation, Cycle 3 High) 3.4 Complete WFC 4381 (HST Medium - Deep Survey - Cycle 3) 3.5 Complete FGS 4892 (Duplicity Among Hyades Stars - Cycle 3 High) 3.6 Complete PC 4644 (The Nuclei of Nearby S0 and Spiral Galaxies -I A PC Snapshot Imaging) 3.7 Complete PC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 3.8 Complete FOC/96 4546 (Imaging of UV Bright Stars in the Galactic Bulge) 3.9 Complete FOS/BL 4662 (Exploring Triton's UV Spectrum: The First 1800 - 2500 A0 UV Survey) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-